### 30 Jan 2024 v1

@@ Abstract
Methods:We combine new 154 and 216 µm dust polarization measurements carried out using the HAWC+ instrument aboard SOFIA with molecular line observations of CN(1–0) and HCO(1–0) from the IRAM 30-meter telescope to determine the magnetic field geometry, and to estimate the plane of the sky magnetic field strength across the NGC 2024 HII region and the surrounding molecular cloud.
Results:The HAWC+ observations show an ordered magnetic field geometry in NGC 2024 that follows the morphology of the expanding HII region and the direction of the main dense filament. The derived plane of the sky magnetic field strength is moderate, ranging from 30 to 80 µG. The strongest magnetic field is found at the northern west edge of the HII region, characterized by the highest gas densities and molecular line widths. In contrast, the weakest field is found toward the main, dense filament in NGC 2024.
Conclusions:We find that the magnetic field has a non negligible influence on the gas stability at the edges of the expanding HII shell (gas impacted by the stellar feedback) and the filament (site of the current star formation).
@@
The question of what controls star formation efficiency in molecular clouds has long been at the center of star formation research. Early studies (Zuckerman & Evans, 1974) showed that if all the gas within dense interstellar clouds were to collapse freely under self-gravity, the star formation rate in the Milky Way would be two orders of magnitude higher than the observed rate of 2 M/year (Robitaille & Whitney, 2010). Theories proposed to explain such a low star formation rate invoke the presence of turbulence or magnetic fields supporting interstellar clouds against gravitational collapse. In some models (Tan et al., 2006), turbulent or magnetic pressure gradients are strong enough to maintain clouds in approximate hydrostatic equilibrium on all spatial scales for tens of free-fall times. In other models, high-density regions are rapidly contracting, converting a large fraction (as high as 40%; Bonnell et al., 2011) of their mass into stars in only a few free-fall times. However, the low-density parts of the cloud, which contain up to 90% of the cloud mass (Battisti & Heyer, 2014), disperse over the scales of free fall time as a result of turbulence generated by galactic shear or by energy input from supernovae explosions (Dobbs et al., 2011; Walch & Naab, 2015), or due to radiative and mechanical stellar feedback from high-mass stars formed early on during the evolution of the cloud (Murray, 2011; Colín et al., 2013; Pabst et al., 2020; Chevance et al., 2022; Suin et al., 2022).
@@
However, it has been difficult to test these star formation models observationally. Despite all the efforts over the past 40 years to characterize interstellar turbulence, we still struggle to understand the different energy sources that contribute to the observed line of sight gas velocity dispersion. The only available way to measure turbulent motions is by observing the gas velocity along the line of sight (Larson, 1981; Hennebelle & Falgarone, 2012), which does not provide a complete picture. As noted by Ballesteros-Paredes et al. (2011), a collapsing cloud will have energetic properties (at least in terms of the total kinetic energy) similar to those of an identical cloud supported against gravity by turbulence. In addition, measurements of magnetic fields are observationally extremely challenging, as they typically involve relatively weak signals, as is the case for Zeeman gas line splitting (Crutcher, 2012) or (sub)millimeter/far-infrared (FIR) dust polarization measurements (see, e.g., Pattle et al., 2019, for a recent review).
@@
Interstellar dust thermal emission is polarized (Hall & Mikesell, 1949; Hiltner, 1949) due to the presence of -fields and non-spherical dust grains in the interstellar medium (ISM). The explanation for this phenomenon is firstly proposed by Davis & Greenstein (1951) as a paramagnetic alignment with the magnetic field. Later on, this process is described by the Radiative Alignment Torques theory (RAT; Hoang & Lazarian, 2014; Andersson et al., 2015, and references therein), in which the minor axis of dust grains is aligned parallel to the direction of the -field. Consequently, the (sub)millimeter dust continuum emission will be polarized perpendicular to the direction of the component of the -field in the plane of the sky (POS).
@@
Regardless of the observational challenges, technological advances in the past decade are now allowing measuring polarized dust emission over increasingly extended regions. The Planck satellite provided an all-sky map of the magnetic field geometry in the diffuse ISM, albeit at a low angular resolution of 5–15. These observations revealed that the galactic -field is intertwined with the filamentary structure of the ISM. In particular, the plane-of-the-sky orientation of the -field and the filamentary structures is correlated with the observed column density, with the field largely parallel to diffuse structures (i.e., striations) while perpendicular to dense filaments (Planck Collaboration et al., 2016a, b). While supported by independent studies (McClure-Griffiths et al., 2006; Goldsmith et al., 2008; Peretto et al., 2012; Clark & Hensley, 2019), this picture is complicated by potential projection effects (Panopoulou et al., 2016) or the presence of stellar feedback (e.g., HII bubble shells, Chapman et al., 2011).
@@
BLASTPol has also observed similar -field morphology at higher angular resolution (2.5, Fissel et al., 2019), which has been interpreted as evidence of the magnetic field dominating the energy balance in the diffuse gas. In contrast, gravity is dominant in the dense, star-forming regions. This interpretation has recently been reinforced by high angular resolution () ground-based observations using the POL-2 instrument on the James Clerk Maxwell Telescope (JCMT; e.g., Pattle et al., 2017; Liu et al., 2019; Wang et al., 2019), which have shown that field lines bend along dense star-forming filaments, possibly as a result of gravitationally-driven flows (Goldsmith et al., 2008; Chapman et al., 2011; Hwang et al., 2023).
@@
A transition from magnetically-dominated to gravitationally-dominated gas requires a redistribution of magnetic flux (e.g., Tritsis et al., 2022). In principle, such a transition occurs before the cores form (Ching et al., 2022), which implies that cores should be generally super-critical. Nevertheless, a change in the gravitational stability of the gas should be accompanied by a corresponding change in its kinematic properties, arguing for the necessity of combining high angular resolution dust polarization and velocity-resolved molecular emission imaging. Such studies have only recently started and cover relatively small isolated fields. Indeed, a recent study (Tang et al., 2019) has investigated the relation between dense gas velocity gradients (as imaged in NH) and the magnetic field direction in massive star-forming infrared dark clouds and concluded that the two are strongly correlated as a result of gravity dragging matter towards the center of the massive ridge. Extending such studies to larger fields is of crucial importance and the High-resolution Airborne Wide-band Camera-Plus (HAWC+, Harper et al., 2018) instrument on the Stratospheric Observatory for Infrared Astronomy (SOFIA) offered exceptional capabilities for mapping the magnetic field geometry, as demonstrated, e.g., by the observations of the OMC-1 region at 53, 89, 154, and 214 µm (Chuss et al., 2019). The capabilities of HAWC+ for large-area magnetic field mapping have further improved with the commissioning of the on-the-fly map (OTFMAP) polarimetric mode for fast wide-field polarimetric imaging.
@@
For estimating the strength of the POS -field, it is critical to know the level of turbulence in the gas because the small-scale turbulence causes local deviations from the mean direction of the magnetic field. The Davis-Chandrasekhar-Fermi method (DCF; Davis & Greenstein, 1951; Chandrasekhar & Fermi, 1953) provides a way of calculating the POS -field strength using information about gas density, turbulence, and changes in the direction of the magnetic field. This method assumes linear geometry and sub-Alfvénic (magnetically-dominated) turbulence. Therefore, it is necessary to obtain information about the turbulence of the gas and dust polarization to use this method. Several variations of the DCF have been developed over the last few decades. The main modification lies in taking into account the number of turbulent cells present along the line of sight and those captured within the beam by including a correction factor Q, which takes a value between 0 and 1 (see Eq. 16 in Ostriker et al., 2001). In another variation, the deviation in the mean direction of the -field is substituted by the ratio between ordered and turbulent magnetic field component (Houde et al., 2009, and references therein). The DCF method assumes that imcompressible motions cause the dispersions of the observed polarization angles, which is not always applicable within the ISM. Skalidis & Tassis (2021) provided an alternative method for deriving magnetic field strength from dust polarization measurements (Skalidis-Tassis method; hereafter ST). In addition to Alfvénic modes, this study accounts for the presence of compressible motions in the gas without discarding the anisotropic nature of the turbulence, providing a physically-motivated approach to estimating the magnetic field strength (Skalidis et al., 2021). In this work, we use the ST method to investigate highly structured regions in the Flame Nebula in the Orion B complex, while also presenting the results from the modified DCF method (e.g., Ostriker et al., 2001; Crutcher, 2004; Lyo et al., 2021) for comparison.
@@
Although OMC-1 is the closest and best-studied high-mass star-forming region, it is not a typical cloud to study the role of the magnetic field in star formation. The reason is that OMC-1 is particularly dense (10 cm) and very active in terms of star formation, affected by strong shocks (including an explosive outflow resulting from the merger of 3 massive protostars, see, e.g., Bally, 2008). These shocks exhibit exceptionally intense UV illumination, with enhancement factors of 10, up to 10 relative to the standard interstellar radiation field (ISRF) (in the vicinity of the Trapezium cluster, Goicoechea et al., 2015, 2019). The presence of a strong radiation field implies that the thermal pressure is high, and the magnetic field plays a limited role in the photon-dominated region (PDR) gas dynamics.
@@
Fortunately, the vast quantity of molecular data already obtained in Orion B makes this region a particularly well-suited location for studying the correlation between changes in the magnetic field geometry and associated gas velocity gradients. The Orion B giant molecular cloud (GMC) hosts NGC 2024, one of the closest high-mass star-forming regions (at a distance of pc; Cao et al., 2023), and is overall more representative of a standard GMC in our galaxy and normal galaxies, with FUV radiation field in the range of (Santa-Maria et al., 2023).
@@
In a recent study, Orkisz et al. (2017) used CO(1–0) and CO(1–0) lines to characterize observationally the ratio of compressive versus solenoidal motions in the turbulent flow and to relate this to the star formation efficiency in various regions of Orion B. Orkisz et al. (2019) accurately analyzed the dynamics of the filamentary network using CO(1–0). Most identified filaments in Orion B are low-density, thermally sub-critical structures, not collapsing to form stars. Only about 1% of the Orion B cloud mass can be found in super-critical, star-forming filaments, consistent with the low overall star formation efficiency of the region (Orkisz et al., 2019).
@@
NGC 2024 (Fig. 1), is located east of Alnitak ( Ori) in the Orion B complex (e.g., Meyer et al., 2008). This region contains a massive and young HII region (age yr, Tremblin et al., 2014, and private communication), deeply embedded in dust, located in the foreground, and extending to the south (e.g., Barnes et al., 1989, see gray contrours showing the network of filaments from Orkisz et al. (2019) in the left panel and labels in the right panel in Fig. 1). The north-south filament in NGC 2024 is super-critical (Orkisz et al., 2019) and the site of ongoing star formation as recently confirmed in the southern part of this filament observed by NOEMA (Shimajiri et al., 2023).
@@
The dust bar observed along the line of sight is visible in the ESO-Vista image (right panel in Fig. 1) as the dark extinction pattern across the image, leading to an apparent cooler dust temperature derived from Herschel observations (Lombardi et al., 2014). The central part of NGC 2024 contains warm dust and gas, heated by the HII region, as well as embedded protostellar objects (Mezger et al., 1988, 1992; Lis et al., 1991) located in the background (light blue points on the right panel in Fig. 1). The young HII region is expanding, strongly impacting its parental cloud, and creating sharp ionization fronts towards the south, which makes it a good example demonstrating how such systems can efficiently exert stellar feedback. The edges of the bubble are seen towards the west and east of the center of NGC 2024 (labeled as Eastern and Western Loop in Fig. 1).
@@ 2.1 Dust polarization measurements using SOFIA HAWC+
Our work employs the dust polarization measurements acquired using HAWC+ on SOFIA in September 2021 for program 09_0154 (PI: D. Lis). Specifically, we observed NGC 2024 at 154 µm (Band D) and 214 µm (Band E) with HAWC+ in polarization mode on flights F779 (2021/09/08), F780 (2021/09/09), F782 (2021/09/11), F783 (2021/09/14), and F784 (2021/09/15). The map obtained at each wavelength comprises ten strips in a weave pattern: five vertical and five horizontal. These strips were designed using the On-The-Fly scan mode of the HAWC+ camera, which covers the requested area using Lissajous patterns on the sky (Harper et al., 2018). Each strip was repeated at least twice.
@@
The polarized intensity has to be corrected for the bias created by the quadratic addition of the noise in the Stokes and maps (Wardle & Kronberg, 1974; Naghizadeh-Khouei & Clarke, 1993). This de-biased polarized intensity (although this is not the only way to do so, e.g., Montier et al., 2015), used in this work, is calculated using
@@
Prior to the data analysis, we filter out low signal-to-noise (S/N) data points. After masking the data, we keep pixels that show S/N ratio in total intensity, in polarization angle uncertainty, and in polarization fraction. Our HAWC+ observations filter out a non-negligible fraction of low-level extended emission. We estimate the amount of missing flux in the SOFIA observations by comparing our data with PACS measurements in Appendix A.1. The data set used in this study was reduced using a larger number (40) of iterations, or rounds, than the default (15) used for the original Level 4 data available on IRSA. A larger number of iterations during data reduction typically improves the recovery of diffuse large-scale emission for scan mode maps. We compare our reduced maps of each Stokes parameter and the polarization angle with those derived from the standard pipeline setup and show them in App. A.1 in Figs. 7 and 8. We find an overall good agreement between these two data sets.

### 07 Feb 2024 v2

@@ Abstract
Methods:We combined new 154 and 216 µm dust polarization measurements carried out using the HAWC+ instrument aboard SOFIA with molecular line observations of CN(1–0) and HCO(1–0) from the IRAM 30-meter telescope to determine the magnetic field geometry, and to estimate the plane of the sky magnetic field strength across the NGC 2024 HII region and the surrounding molecular cloud.
Results:The HAWC+ observations show an ordered magnetic field geometry in NGC 2024 that follows the morphology of the expanding HII region and the direction of the main dense filament. The derived plane of the sky magnetic field strength is moderate, ranging from 30 to 80 µG. The strongest magnetic field is found at the east edge of the HII region, characterized by the highest gas densities and molecular line widths. In contrast, the weakest field is found toward the main, dense filament in NGC 2024.
Conclusions:We find that the magnetic field has a non-negligible influence on the gas stability at the edges of the expanding HII shell (gas impacted by stellar feedback) and the filament (site of current star formation).
@@
The question of what controls the star formation efficiency in molecular clouds has long been at the center of star formation research. Early studies (Zuckerman & Evans, 1974) showed that if all the gas within dense interstellar clouds were to collapse freely under self-gravity, the star formation rate in the Milky Way would be two orders of magnitude higher than the observed rate of 2 M/year (Robitaille & Whitney, 2010). Theories proposed to explain such a low star formation rate invoke the presence of turbulence or magnetic fields supporting interstellar clouds against gravitational collapse. In some models (Tan et al., 2006), turbulent or magnetic pressure gradients are strong enough to maintain clouds in approximate hydrostatic equilibrium on all spatial scales for tens of free-fall times. In other models, high-density regions are rapidly contracting, converting a large fraction (as high as 40%; Bonnell et al., 2011) of their mass into stars in only a few free-fall times. However, the low-density parts of the cloud, which contain up to 90% of the cloud mass (Battisti & Heyer, 2014), disperse over the scales of free fall time as a result of turbulence generated by galactic shear or by energy input from supernova explosions (Dobbs et al., 2011; Walch & Naab, 2015), or due to radiative and mechanical stellar feedback from high-mass stars that formed early on during the evolution of the cloud (Murray, 2011; Colín et al., 2013; Pabst et al., 2020; Chevance et al., 2022; Suin et al., 2022).
@@
However, it has been difficult to test these star formation models observationally. Despite all of the effort made over the past 40 years to characterize interstellar turbulence, we still struggle to understand the different energy sources that contribute to the observed line-of-sight gas velocity dispersion. The only available way to measure turbulent motions is by observing the gas velocity along the line of sight (Larson, 1981; Hennebelle & Falgarone, 2012), which does not provide a complete picture. As noted by Ballesteros-Paredes et al. (2011), a collapsing cloud has energetic properties (at least in terms of the total kinetic energy) similar to those of an identical cloud supported against gravity by turbulence. In addition, measurements of magnetic fields are observationally extremely challenging, as they typically involve relatively weak signals, as is the case for Zeeman gas line splitting (Crutcher, 2012) or (sub)millimeter/far-infrared (FIR) dust polarization measurements (see, Pattle et al., 2019, for a recent review).
@@
Interstellar dust thermal emission is polarized (Hall & Mikesell, 1949; Hiltner, 1949) due to the presence of fields and non-spherical dust grains in the interstellar medium (ISM). The explanation for this phenomenon was firstly proposed by Davis & Greenstein (1951) as a paramagnetic alignment with the magnetic field. Later on, this process was described by the radiative alignment torques (RAT) theory (Hoang & Lazarian, 2014; Andersson et al., 2015, and references therein), in which the minor axis of dust grains is aligned parallel to the direction of the field. Consequently, the (sub)millimeter dust continuum emission is polarized perpendicular to the direction of the component of the field in the plane of the sky (POS).
@@
Regardless of the observational challenges, technological advances in the past decade are now allowing for polarized dust emission to be measured over increasingly extended regions. The Planck satellite provided an all-sky map of the magnetic field geometry in the diffuse ISM, albeit at a low angular resolution of 5–15. These observations revealed that the galactic field is intertwined with the filamentary structure of the ISM. In particular, the POS orientation of the field and the filamentary structures is correlated with the observed column density, with the field largely parallel to diffuse structures (striations) while perpendicular to dense filaments (Planck Collaboration et al., 2016a, b). While supported by independent studies (McClure-Griffiths et al., 2006; Goldsmith et al., 2008; Peretto et al., 2012; Clark & Hensley, 2019), this picture is complicated by potential projection effects (Panopoulou et al., 2016) or the presence of stellar feedback (HII bubble shells, Chapman et al., 2011).
@@
The Balloon-borne Large Aperture Submillimeter Telescope for Polarimetry (BLASTPol) has also observed a similar -field morphology at a higher angular resolution (2.5, Fissel et al., 2019), which has been interpreted as evidence of the magnetic field dominating the energy balance in the diffuse gas. In contrast, gravity is dominant in the dense, star-forming regions. This interpretation has recently been reinforced by high angular resolution () ground-based observations using the POL-2 instrument on the James Clerk Maxwell Telescope (JCMT; Pattle et al., 2017; Liu et al., 2019; Wang et al., 2019), which have shown that field lines bend along dense star-forming filaments, possibly as a result of gravitationally driven flows (Goldsmith et al., 2008; Chapman et al., 2011; Hwang et al., 2023).
@@
A transition from magnetically dominated to gravitationally dominated gas requires a redistribution of magnetic flux (Tritsis et al., 2022). In principle, such a transition occurs before the cores form (Ching et al., 2022), which implies that cores should be generally super-critical. Nevertheless, a change in the gravitational stability of the gas should be accompanied by a corresponding change in its kinematic properties, arguing for the necessity of combining high angular resolution dust polarization and velocity-resolved molecular emission imaging. Such studies have only recently started and cover relatively small isolated fields. A recent study (Tang et al., 2019) has indeed investigated the relation between dense gas velocity gradients (as imaged in NH) and the magnetic field direction in massive star-forming infrared dark clouds and concluded that the two are strongly correlated as a result of gravity dragging matter toward the center of the massive ridge. Extending such studies to larger fields is of crucial importance and the High-resolution Airborne Wide-band Camera-Plus (HAWC+, Harper et al., 2018) instrument on the Stratospheric Observatory for Infrared Astronomy (SOFIA) offered exceptional capabilities for mapping the magnetic field geometry, as demonstrated, for example, by the observations of the OMC-1 region at 53, 89, 154, and 214 µm (Chuss et al., 2019). The capabilities of HAWC+ for large-area magnetic field mapping have further improved with the commissioning of the on-the-fly map (OTFMAP) polarimetric mode for fast wide-field polarimetric imaging.
@@
When estimating the strength of the POS field, it is critical to know the level of turbulence in the gas because the small-scale turbulence causes local deviations from the mean direction of the magnetic field. The Davis-Chandrasekhar-Fermi (DCF) method (Davis & Greenstein, 1951; Chandrasekhar & Fermi, 1953) provides a way of calculating the POS -field strength using information about gas density, turbulence, and changes in the direction of the magnetic field. This method assumes linear geometry and sub-Alfvénic (magnetically dominated) turbulence. Therefore, it is necessary to obtain information about the turbulence of the gas and dust polarization to use this method. Several variations of the DCF have been developed over the last few decades. The main modification lies in taking into account the number of turbulent cells present along the line of sight and those captured within the beam by including a correction factor Q, which takes a value between zero and one (see Eq. 16 in Ostriker et al., 2001). In another variation, the deviation in the mean direction of the field is substituted by the ratio between the ordered and turbulent magnetic field component (Houde et al., 2009, and references therein). The DCF method assumes that imcompressible motions cause the dispersions of the observed polarization angles, which is not always applicable within the ISM. Skalidis & Tassis (2021) provided an alternative method for deriving the magnetic field strength from dust polarization measurements – the Skalidis-Tassis (ST) method. In addition to Alfvénic modes, this study accounts for the presence of compressible motions in the gas without discarding the anisotropic nature of the turbulence, providing a physically motivated approach to estimating the magnetic field strength (Skalidis et al., 2021). For this work, we used the ST method to investigate highly structured regions in the Flame nebula in the Orion B complex, while also presenting the results from the modified DCF method (for example, Ostriker et al., 2001; Crutcher, 2004; Lyo et al., 2021) for comparison.
@@
Although OMC-1 is the closest and best-studied high-mass star-forming region, it is not a typical cloud to study the role of the magnetic field in star formation. The reason is that OMC-1 is particularly dense (10 cm) and very active in terms of star formation, affected by strong shocks (including an explosive outflow resulting from the merger of three massive protostars, see, Bally, 2008). These shocks exhibit exceptionally intense UV illumination, with enhancement factors of 10, up to 10 relative to the standard interstellar radiation field (ISRF) (in the vicinity of the Trapezium cluster, Goicoechea et al., 2015, 2019). The presence of a strong radiation field implies that the thermal pressure is high, and the magnetic field plays a limited role in the photon-dominated region (PDR) gas dynamics.
@@
Fortunately, the vast quantity of molecular data already obtained in Orion B makes this region a particularly well-suited location for studying the correlation between changes in the magnetic field geometry and associated gas velocity gradients. The Orion B giant molecular cloud (GMC) hosts NGC 2024, one of the closest high-mass star-forming regions (at a distance of pc; Cao et al., 2023), and it is overall more representative of a standard GMC in our galaxy and normal galaxies, with the far-ultraviolet (FUV) radiation field in the range of (Santa-Maria et al., 2023).
@@
In a recent study, Orkisz et al. (2017) used CO(1–0) and CO(1–0) lines to characterize observationally the ratio of compressive versus solenoidal motions in the turbulent flow and to relate this to the star formation efficiency in various regions of Orion B. Orkisz et al. (2019) accurately analyzed the dynamics of the filamentary network using CO(1–0). Most identified filaments in Orion B are low-density, thermally subcritical structures, not collapsing to form stars. Only about 1% of the Orion B cloud mass can be found in super-critical, star-forming filaments, consistent with the low overall star formation efficiency of the region (Orkisz et al., 2019).
@@
NGC 2024 (Fig. 1) is located east of Alnitak ( Ori) in the Orion B complex (Meyer et al., 2008). This region contains a massive and young HII region (age yr, Tremblin et al., 2014, and private communication), deeply embedded in dust, located in the foreground, and extending to the south (Barnes et al., 1989, see gray contrours showing the network of filaments from Orkisz et al. (2019) in the left panel and labels in the right panel in Fig. 1). The north-south filament in NGC 2024 is super-critical (Orkisz et al., 2019) and the site of ongoing star formation as recently confirmed in the southern part of this filament observed by NOEMA (Shimajiri et al., 2023).
@@
The dust bar observed along the line of sight is visible in the ESO-Vista image (right panel in Fig. 1) as the dark extinction pattern across the image, leading to an apparent cooler dust temperature derived from Herschel observations (Lombardi et al., 2014). The central part of NGC 2024 contains warm dust and gas, heated by the HII region, as well as embedded protostellar objects (Mezger et al., 1988, 1992; Lis et al., 1991) located in the background (light blue points in the right panel in Fig. 1). The young HII region is expanding, strongly impacting its parental cloud, and creating sharp ionization fronts toward the south, which makes it a good example demonstrating how such systems can efficiently exert stellar feedback. The edges of the bubble are seen toward the west and east of the center of NGC 2024 (labeled as Eastern and Western Loop in Fig. 1).
@@ 2.1 Dust polarization measurements using SOFIA HAWC+
Our work employs the dust polarization measurements acquired using HAWC+ on SOFIA in September 2021 for program 09_0015 (PI: D. Lis). Specifically, we observed NGC 2024 at 154 µm (Band D) and 214 µm (Band E) with HAWC+ in polarization mode on flights F779 (8 September 2021), F780 (9 September 2021), F782 (11 September 2021), F783 (14 September 2021), and F784 (15 September 2021). The map obtained at each wavelength comprises ten strips in a weave pattern: five vertical and five horizontal. These strips were designed using the On-The-Fly scan mode of the HAWC+ camera, which covers the requested area using Lissajous patterns on the sky (Harper et al., 2018). Each strip was repeated at least twice.
@@
The polarized intensity has to be corrected for the bias created by the quadratic addition of the noise in the Stokes and maps (Wardle & Kronberg, 1974; Naghizadeh-Khouei & Clarke, 1993). This de-biased polarized intensity (although this is not the only way to do so, Montier et al., 2015), used in this work, is calculated using
@@
Prior to the data analysis, we filter out low signal-to-noise ratio (S/N) data points. After masking the data, we keep pixels that show an S/N in total intensity, in polarization angle uncertainty, and in polarization fraction. Our HAWC+ observations filter out a non-negligible fraction of low-level extended emission. We estimate the amount of missing flux in the SOFIA observations by comparing our data with PACS measurements in Appendix A.1. The data set used in this study was reduced using a larger number (40) of iterations, or rounds, than the default (15) used for the original Level 4 data available on IRSA. A larger number of iterations during data reduction typically improves the recovery of diffuse large-scale emission for scan mode maps. We compare our reduced maps of each Stokes parameter and the polarization angle with those derived from the standard pipeline setup and show them in App. A.1 in Figs. 7 and 8. We find an overall good agreement between these two data sets.